# VerifyBench: Benchmarking Reference-based Reward Systems for Large Language Models

### v2

@@ 4.1 Overall Performance
Skywork/Skywork-Reward-Llama-3.1-8B 72.29 93.33 60.80 64.80 59.60 68.80 63.48
internlm/internlm2-20b-reward 72.06 92.16 65.60 64.80 61.20 70.00 65.40
Ray2333/GRM-llama3-8B-sftreg 71.33 88.50 64.80 58.40 58.80 67.60 62.40
internlm/internlm2-7b-reward 72.42 90.02 73.20 68.00 66.80 70.40 69.60
@@
Qwen/Qwen2.5-Math-RM-72B 76.28 82.11 83.60 79.20 73.60 75.60 78.00
Qwen/Qwen2-Math-RM-72B 62.61 75.54 79.20 78.40 73.20 72.80 75.90
4.2 Reference-answers Play an Important Role in Verification
The benchmark proposed in this work fundamentally differs from existing reward benchmarks by explicitly incorporating reference answers, thereby aligning more closely with the training setups of contemporary reasoning LLMs. To isolate the impact of reference answers on verification performance, we conduct an ablation study in which models are evaluated without reference answers provided in the prompt; the prompt format used is detailed in Appendix C.3.
Experimental results, summarized in Table 3, reveal a performance degradation of approximately 5–18% when reference answers are excluded. These findings underscore the crucial role of reference answers in reasoning-oriented RL, suggesting they provide a more reliable and informative supervision signal during reward modeling.
4.3 Performance of Reference-free Reward Models
Our experimental results show that VerifyBench introduces a level of challenge comparable to existing reward benchmarks, with the absence of reference answers. Reference-free reward models achieve sub-80% accuracy on VerifyBench, highlighting its difficulty. Furthermore, domain-specific reward models exhibit inferior performance on general reward benchmarks compared to VerifyBench, validating the benchmark’s design objectives.
@@ 5.1 Error Analysis
To gain deeper insights from VerifyBench, we introduce a more fine-grained taxonomy for each answer type and analyze model performance across these subcategories. This detailed analysis helps identify specific reasoning tasks or answer formats where models are particularly error-prone. We subdivide the Numeric Values category into 8 subtypes, Expressions into 5 subtypes, Multi-choice into 3 subtypes, and String into 2 subtypes. Table 5 presents the comparative performance of different models across these detailed categories.
  • Expressions: Algebraic formulas and equations;
  • Multi-choice: Multi-answer choice problems;
  • String: Strings requiring semantic consistency verification.
@@ 5.2 Correlation Analysis
In our experiments, we applied rejection sampling to implement reference-based reward systems. For each question in the GSM8K and MATH training sets, we generated 64 candidate completions using Qwen2.5-Math-7B-Instruct (Yang et al., 2024) with a sampling temperature of 0.7. These responses were subsequently filtered by three verifier models with varying performance levels on VerifyBench: Llama-3.1-8B-Instruct, Qwen3-4B, and Qwen3-1.7B. Only completions consistently verified as correct were retained to form the SFT training data. We conducted independent SFT training runs accordingly, with full hyperparameter configurations provided in the Appendix D.
@@ 6 Conclusion
Ethical Considerations
Appendix A Related Works
A.1 Reward Systems for Reinforcement Learning
Early reward models (RMs) (Christiano et al., 2023; Stiennon et al., 2022; Ouyang et al., 2022), trained to predict human preference rankings, typically treat the entire response as the evaluation unit. However, such outcome-level RMs lack insight into intermediate reasoning steps, making step-level error correction infeasible (Xu et al., 2025b). To address this limitation, process-level RMs (Lightman et al., 2023; Setlur et al., 2024) have been introduced to assign scores at each reasoning step, thereby providing stepwise feedback. Despite their effectiveness, process-level RMs require extensive manual step-level annotations, resulting in exponential increases in data collection costs and training complexity (Khalifa et al., 2025).
@@ A.1 Reward Systems for Reinforcement Learning
More recently, DeepSeek-GRM (Liu et al., 2025) and ThinkPRM (Khalifa et al., 2025) have explored integrating reasoning capabilities into RMs by developing generative reward models (GRMs). GRMs reformulate the scoring task as a token-generation problem: before outputting a numerical score, the model first generates a chain-of-thought (CoT) (Wei et al., 2023) that explicates its evaluation criteria and rationale. This approach not only bridges the interpretability gap between black-box discriminative models and brittle rule-based systems but also substantially enhances test-time scaling capabilities.
A.2 Evaluation of Reward Systems
@@
Reward system benchmarks can be further categorized into outcome-level (Liu et al., 2024; Lambert et al., 2024) and process-level (Lightman et al., 2023; Zheng et al., 2024) suites. In constructing these benchmarks, researchers generate multiple responses to the same prompt by varying model architectures or hyperparameters. During the manual annotation phase, outcome-level benchmarks require annotators to compare or assign multi-point scores to complete responses, emphasizing overall preference. In contrast, process-level benchmarks provide fine-grained gold verdicts by requiring step-by-step correctness labels for each reasoning step.
Beyond benchmark-based evaluation, practical applications of reward systems serve as another common assessment method. In the Best-of-N (BoN) paradigm, WebGPT (Nakano et al., 2022) introduced using a reward model to score $N$ candidate answers and select the top-ranked response. Subsequent work has framed reward models as downstream rankers—for example, Self-Consistency in chain-of-thought models (Wang et al., 2023), where the reward model identifies the most coherent solution among candidates. Unlike Best-of-N, rejection sampling fine-tuning (RFT) (Zelikman et al., 2024; Xiong et al., 2025) samples multiple trajectories from the current policy, scores them using a reward model, and retains only the highest-scoring examples as silver supervision for further fine-tuning. This approach has proven particularly effective at bootstrapping reasoning capabilities without requiring full preference-learning pipelines.
Appendix B Data Source
@@
Table 6: LLMs used in this paper.
Appendix C Prompts
C.1 Prompt for Answer Type Classification
@@
We present the prompt we used to generate answer types in Figure 4.
C.2 Prompt for LLM-as-a-judge
@@
We present the prompt we used in LLM-as-a-judge evaluation with a reference answer in Figure 5.
C.3 Prompt for LLM-as-a-judge without Reference
@@
We present the prompt we used in LLM-as-a-judge evaluation with a reference answer in Figure 6.
Appendix D Experimental Details

### v3

@@ 4.3 Performance of Reference-free Reward Models
Skywork-Reward-Llama-3.1-8B 72.29 93.33 60.80 64.80 59.60 68.80 63.48
internlm2-20b-reward 72.06 92.16 65.60 64.80 61.20 70.00 65.40
GRM-llama3-8B-sftreg 71.33 88.50 64.80 58.40 58.80 67.60 62.40
internlm2-7b-reward 72.42 90.02 73.20 68.00 66.80 70.40 69.60
@@
Qwen2.5-Math-RM-72B 76.28 82.11 83.60 79.20 73.60 75.60 78.00
Qwen2-Math-RM-72B 62.61 75.54 79.20 78.40 73.20 72.80 75.90
@@ 5.1 Error Analysis
To gain deeper insights from VerifyBench, we introduce a more fine-grained taxonomy for each answer type and analyze model performance across these subcategories. This detailed analysis helps identify specific reasoning tasks or answer formats where models are particularly error-prone. We subdivide the Numeric Values category into 8 subtypes, Expressions into 5 subtypes, Multi-choice into 3 subtypes, and String into 2 subtypes. Table 5 presents the comparative performance of different models across these detailed categories. The complete taxonomy of VerifyBench and VerifyBench-Hard is provided in Appendix G.
  • Expressions: Formulas and equations;
  • Multi-choice: Multiple-choice problems;
  • String: Strings requiring semantic consistency verification.
@@
Figure 3: The performance(%) of RL across different LLM judges which have various performance on VerifyBench.
@@ 5.2 Correlation Analysis
We conducted practical RL training on Qwen2.5-3B-Instruct (Qwen et al., 2025), using LLMs with varying performance on VerifyBench as LLM judgers to serve as the reward system. We then compared their improvements on downstream tasks. Specifically, we adopted Llama-3.2-3B-Instruct (Grattafiori et al., 2024), Qwen3-1.7, Qwen3-4B (Yang et al., 2025), and gpt-oss-20b (OpenAI et al., 2025) as verifiers, employed GRPO (Shao et al., 2024) as the RL algorithm, and trained on DeepMath-103K. The accuracy changes on MATH500 (Hendrycks et al., 2021), AMC23 (mathai, 2023), and MinervaMath (Lewkowycz et al., 2022) are shown in Figure 3. Meanwhile, we provide a additional correlation analysis based on Rejection Sampling Fine-tuning and all the experimental details in Appendix H.
From our experimental results, we observe that models performing well on VerifyBench, when used as verifiers in RL training, lead to greater improvements on related tasks. For example, employing gpt-oss-20b and Qwen3-4B yields significant gains, whereas using models with weaker VerifyBench performance as verifiers may even degrade the model’s capability. This is because such models introduce substantial bias in reward assignment, resulting in inaccurate supervision. Our experiments demonstrate that VerifyBench provides effective guidance in RL training scenarios, highlighting the practical utility of our benchmarks.
@@ Appendix A Limitations
Appendix B Related Works
B.1 Reward Systems for Reinforcement Learning
Early reward models (RMs) (Christiano et al., 2017; Stiennon et al., 2020; Ouyang et al., 2022), trained to predict human preference rankings, typically treat the entire response as the evaluation unit. However, such outcome-level RMs lack insight into intermediate reasoning steps, making step-level error correction infeasible (Xu et al., 2025b). To address this limitation, process-level RMs (Lightman et al., 2023; Setlur et al., 2024) have been introduced to assign scores at each reasoning step, thereby providing stepwise feedback. Despite their effectiveness, process-level RMs require extensive manual step-level annotations, resulting in exponential increases in data collection costs and training complexity (Khalifa et al., 2025).
@@ B.1 Reward Systems for Reinforcement Learning
More recently, DeepSeek-GRM (Liu et al., 2025b) and ThinkPRM (Khalifa et al., 2025) have explored integrating reasoning capabilities into RMs by developing generative reward models (GRMs). GRMs reformulate the scoring task as a token-generation problem: before outputting a numerical score, the model first generates a chain-of-thought (CoT) (Wei et al., 2022) that explicates its evaluation criteria and rationale. This approach not only bridges the interpretability gap between black-box discriminative models and brittle rule-based systems but also substantially enhances test-time scaling capabilities.
With the advancement of Reinforcement Learning with Verifiable Reward (RLVR), using reference answers to verify correctness on reasoning tasks has become the mainstream approach for RL training of reasoning models. Most existing work employs rule-based verifiers to assess the consistency between rollouts and reference answers (DeepSeek-AI et al., 2025a; He et al., 2025), thereby providing rewards in RL. These rule-based functions typically rely on extensive human-crafted heuristic rules and often lack verification robustness (Hong et al., 2025). Consequently, recent work has incorporated reference answers into reward models, leveraging the generalization capabilities of model-based verifiers to improve verification accuracy (Chen et al., 2025; Liu et al., 2025a; Su et al., 2025).
B.2 Evaluation of Reward Systems
@@
Reward system benchmarks can be further categorized into outcome-level (Liu et al., 2024; Lambert et al., 2025) and process-level (Lightman et al., 2023; Zheng et al., 2025) suites. In constructing these benchmarks, researchers generate multiple responses to the same prompt by varying model architectures or hyperparameters. During the manual annotation phase, outcome-level benchmarks require annotators to compare or assign multi-point scores to complete responses, emphasizing overall preference. In contrast, process-level benchmarks provide fine-grained gold verdicts by requiring step-by-step correctness labels for each reasoning step.
Beyond benchmark-based evaluation, practical applications of reward systems serve as another common assessment method. In the Best-of-N (BoN) paradigm, WebGPT (Nakano et al., 2022) introduced using a reward model to score $N$ candidate answers and select the top-ranked response. Subsequent work has framed reward models as downstream rankers, for example, Self-Consistency in chain-of-thought models (Wang et al., 2022), where the reward model identifies the most coherent solution among candidates. Unlike Best-of-N, rejection sampling fine-tuning (RFT) (Zelikman et al., 2024; Xiong et al., 2025) samples multiple trajectories from the current policy, scores them using a reward model, and retains only the highest-scoring examples as silver supervision for further fine-tuning. This approach has proven particularly effective at bootstrapping reasoning capabilities without requiring full preference-learning pipelines.
During the course of this work, we also discovered several concurrent works that constructed evaluation benchmarks for reward systems based on reference answers. Specifically, Liu et al. (2025a) proposed VerifierBench, which employs a similar construction approach to our VerifyBench (easy version). However, compared to our work, they utilized fewer datasets to generate evaluation samples, using 14 data sources while we employed 41 data sources. Additionally, we adopted a more fine-grained classification system for annotation. Furthermore, we propose a disagreement-based sample selection strategy to construct VerifyBench-Hard, which includes samples that are difficult for model-based verifiers to evaluate, thereby enhancing the benchmark’s difficulty. Our benchmarks also incorporate outputs from reasoning models, aligning with the mainstream forms of contemporary model outputs and improving the benchmarks’ coverage.
Appendix C Data Source
@@
The datasets we used mainly fall into three categories: general_reasoning, logic_reasoning, and math_reasoning. Among them, general_reasoning also includes subsets such as commonsense_reasoning. In our selection, general_reasoning consists of 4 datasets, logic_reasoning includes 21 datasets, and math_reasoning comprises 16 datasets, among which one is collected by ourselves. All of the above datasets provide original reference answers. In our dataset selection, we do not differentiate between the training and test splits of the original datasets; instead, we only use their questions and reference answers.
Domain Source License # of Questions
general_reasoning BBH (Suzgun et al., 2023) MIT 4520
BBEH (Kazemi et al., 2025) Apache 2.0 6511
MMLU_pro (Wang et al., 2024) Apache 2.0 2000
natural_reasoning (Yuan et al., 2025) CC-BY-NC 4.0 10000
logic_reasoning AbductionRules (Young et al., 2022) MIT 1000
anlg (Bhagavatula et al., 2019) / 1000
anli (Nie et al., 2020) CC-BY-NC 4.0 1000
ARLSAT (Zhong et al., 2021) MIT 230
bAbI15 (Weston et al., 2015) / 1000
bAbI16 (Weston et al., 2015) / 1000
BoardgameQA (Kazemi et al., 2023) $C$-BY-4.0 1000
clutrr (Sinha et al., 2019) CC-BY-NC 4.0 1000
FOLIO (Han et al., 2024) $C$-BY-SA-4.0 134
hellaswag (Zellers et al., 2019) MIT 1000
logicbenchBQA (Parmar et al., 2024) MIT 1000
logicbenchMCQA (Parmar et al., 2024) MIT 1000
LogiQA (Liu et al., 2020) / 1000
MultiLogiEval (Patel et al., 2024) MIT 1000
NeuLRabductive (Xu et al., 2025a) / 1000
NeuLRdeductive (Xu et al., 2025a) / 1000
NeuLRinductive (Xu et al., 2025a) / 1000
ProntoQA (Saparov & He, 2022) Apache 2.0 500
ProofWriter (Tafjord et al., 2021) / 1000
ReClor (Yu et al., 2019) / 500
tablebench (Wu et al., 2025) Apache 2.0 886
math_reasoning AIME24 MIT 30
AIME25 MIT 30
asdiv-a (Miao et al., 2020) CC-BY-NC 4.0 1218
Math Odyssey (Fang et al., 2025) MIT 389
GPQA_diamond (Rein et al., 2024) MIT 198
gsm8k (Cobbe et al., 2021) MIT 1319
math401 (Yuan et al., 2023) / 392
mathematics (Saxton et al., 2018) Apache 2.0 3360
MATH(Hendrycks et al., 2021) MIT 5000
OlympiadBench-EN (He et al., 2024) MIT 675
SVAMP (Patel et al., 2021) MIT 1000
NuminaMath-CoT (Li et al., 2024) Apache 2.0 20000
orca-math-word-problems (Mitra et al., 2024) MIT 10000
ArtOfProblemSolving self-curated 7997
stackmathqa (Zhang, Yifan, 2024) $C$-BY-4.0 10000
DeepMath-103K-RL (He et al., 2025) MIT 20000
Appendix D LLM Usage For Benchmark Construction
We list all the LLMs we used to generate completions for curated question in Table 7. We carefully selected 22 LLMs from 10 families of both open-source and closed-source models to generate completions. Each model produced one completion for each question, which was collected as candidate completions for subsequent processing.
For all models, we strictly adhered to their licenses and ensured that the generated data would not be used for any model training, but solely for evaluation purposes. For open-source models, we employed the vLLM (Kwon et al., 2023) framework for generation. The generation hyper-parameters were set to temperature=0.7 and top_p=0.95, with all other parameters kept at their default values.
Table 7: All LLMs used to generate completions in this paper. We employed multiple open-source and closed-source models of varying scales to generate completions, thereby enhancing the robustness of our constructed benchmark.
@@ Appendix D LLM Usage For Benchmark Construction
Appendix E Prompts
E.1 Prompt for Answer Type Classification
@@
We present the prompt template we used to generate answer types in Figure 4. {question} and {answer} are placeholders that will be replaced with actual question and answer content.
E.2 Prompt for LLM-as-a-judge
@@
We present the prompt template we used in LLM-as-a-judge evaluation with a reference answer in Figure 5. {question}, {answer}, and {completion} are placeholders that will be replaced with actual question, answer, and completion content.
E.3 Prompt for LLM-as-a-judge without Reference
@@
We present the prompt template we used in LLM-as-a-judge evaluation with a reference answer in Figure 6. {question} and {completion} are placeholders that will be replaced with actual question and completion content.
Figure 4: Prompt for answer type classification. {question} and {answer} are placeholders that will be replaced with actual question and answer content.
Figure 5: Prompt for LLM-as-a-judge evaluation. {question}, {answer}, and {completion} are placeholders that will be replaced with actual question, answer, and completion content.
Figure 6: Prompt for LLM-as-a-judge evaluation without reference answers. {question} and {completion} are placeholders that will be replaced with actual question and completion content.
Appendix F Details of Human Annotation
We present comprehensive details regarding our human annotation process: (1) the composition of our annotation team, including the knowledge backgrounds of annotators, and (2) the annotation interface design (for annotators and for meta-annotators), encompassing annotation instructions and user interface specifications.
F.1 Annotators
Our annotation team consists of 6 members, including 4 annotators and 2 meta-annotators. The meta-annotators are responsible for resolving disagreements that arise during the annotation process.
Among the four annotators, one is an undergraduate student and the other three are master’s students, all of whom have engaged in research on LLM reasoning for over three months. The two meta-annotators are current PhD students who have conducted research on LLM reasoning for more than one year. Among the two meta-annotators, one was responsible for resolving disagreements between general reasoning and math reasoning categories, while the other handled conflicts between general reasoning and logic reasoning categories.
Our primary annotator team consists of the following members:
  • Annotator febd: Undergraduate student with 5 months of experience in LLM reasoning research.
  • Annotator cfaa: Master’s student with 6 months of experience in LLM reasoning research.
  • Annotator ebed: Master’s student with 6 months of experience in LLM reasoning research.
  • Annotator ffce: Master’s degree holder with 1 year of experience in LLM reasoning research.
Our meta-annotator team consists of the following members:
  • Meta-annotator ebed: Doctoral student with 1 year of experience in LLM reasoning research, specializing in reasoning and logical reasoning domains.
  • Meta-annotator feea: Doctoral student with 1 year of experience in LLM reasoning research, specializing in reasoning and mathematical reasoning domains.
F.2 Instruction for Annotators
Before initiating the manual annotation process, we conducted a 60-minute video training session for all annotators. After the training, we also provided written instructions detailing the annotation procedure, which were distributed to all annotators. In addition, these instructions were made accessible on the annotation interface itself, enabling annotators to consult them in real time. The following is the instruction text we provided.
Instruction for Annotators ### What should I do **Step 1:** Check the answer type. The page has been pre-annotated, please check whether this annotation matches the actual situation. Please select the most appropriate option from the dropdown list, where option E is not suitable for VerifyBench. If you select option E, you can skip step 2 and directly click the Next Button.
**Step 2:** Check answer correctness. A total of 4 corresponding answers are displayed, each answer has been pre-annotated for correctness. Please check whether this annotation matches the actual situation, and if it doesn’t match, please make the appropriate corrections.
**Step 3:** Click the Next Button! You have successfully completed the annotation for this case. Thanks for your contribution!
F.3 UI Design for Human Annotation Process
We developed a web-based annotation interface using Streamlit to facilitate human labeling. All annotations were conducted through the web interface, with different annotators distinguished by unique user parameters. We constructed two distinct interfaces: one for primary annotators and another for meta annotators. Primary annotators directly labeled the data, while meta annotators resolved disagreements between primary annotators through our double-check verification mechanism and made final labeling decisions based on the annotations from both primary annotators.
F.3.1 Page for Primary Annotators
We designed an annotation interface for annotators that displays the current annotation progress. By clicking on interface elements, annotators can view the complete data in JSON format to make informed judgments. The interface presents one question, one reference answer, and four corresponding completions at a time. Annotators label the answer category using dropdown components and mark the correctness of the four completions using checkboxes. Our annotation interface supports LaTeX formula rendering, enabling annotators to perform mathematical problem annotation in a more readable format. For VerifyBench annotation, the interface displays pre-annotation results, while for VerifyBench-Hard annotation, no pre-annotation results are shown. The interface incorporates an invalid button to enable annotators to quickly mark samples unsuitable for VerifyBench. Any sample marked as invalid by any annotator is excluded from subsequent annotation processes. Additionally, a skip button is implemented to allow annotators to bypass samples they cannot evaluate with confidence. Figure 7 shows the actual annotation interface for primary annotators.
Figure 7: Annotation interface for primary annotators, one sample per page for annotation.
F.3.2 Page for Meta-Annotators
The interface designed for meta-annotators differs slightly from that of primary annotators. The key distinction lies in the meta-annotators’ ability to view detailed annotations from primary annotators, with disagreements between annotators clearly displayed on the page to facilitate final judgment based on the primary annotation information. Furthermore, we designed domain-based filtering element for meta-annotators, enabling them to select domains aligned with their knowledge background and expertise. Other page logic remains identical to the primary annotators’ interface. Figure 8 illustrates the actual annotation interface for meta-annotators.
Figure 8: Annotation interface for meta-annotators, one sample per page for annotation.
Appendix G Full Taxonomy of VerifyBench and VerifyBench-Hard
During human annotation, we adopted a finer-grained labeling scheme. Beyond the four main categories mentioned in the text (Numeric Values, Expressions, Multi-choice, and String), we further annotated more detailed subcategories within each type for analysis. Specifically, Numeric Values includes 8 subcategories, Expressions includes 5 subcategories, Multi-choice includes 3 subcategories, and String includes 2 subcategories. The descriptions of all subcategories are presented in Table 8.
Table 8: Full taxonomy and their description of VerifyBench and VerifyBench-Hard
Label Type Sub-Label Sub-Type Description
A Numeric Values A1 Integers(whole numbers), e.g. $2$, 351351, etc.
A2 Numerical values containing constants, e.g. $2\pi$, etc.
A3 Fractions/Float numbers/Proportions, e.g. $\frac{1}{2}$, $3.123$, $1:2$, etc.
A4 Numerical values containing radicals, e.g. $\sqrt{3}$
A5 Complex numbers, e.g. $3+4i$
A6 Angles, e.g. $120^{\circ}$
A7 Non-decimal numbers (not base 10 numeric values), e.g. $1234_{8}$
A8 Multiple values with no order between them, e.g. 1,2,11,-2,1 and $2$, etc.
B Expressions B1 Algebraic formulas, e.g. $ax+b$, etc.
B2 Equations or systems of equations, e.g. $ax+b=0$, etc.
B3 Intervals, including concatenation and intersection of intervals, e.g. $(1,2]$ $(1,2)$ $\cup(3,4)$, etc.
B4 A set, a collection of elements, e.g. $\{1,2,3\}$, $(1,2,3)$, etc.
B5 A matrix or vector, e.g. $\begin{pmatrix}-7\\ 16\\ 5\end{pmatrix}$, $(1,2)$ etc.
C Multi-choice C1 Single-choice questions, choose one correct option from multiple choices.
C2 Multiple choice, select multiple correct choices from multiple options.
C3 Finite state selection, choose one correct state from a finite number of states, e.g. True/False, yes/no, valid/invalid, etc.
D String D1 Specific natural language representation, more concerned with whether a particular word or expression is mentioned correctly or not
D2 Semantically consistent natural language representations, where the answer is difficult to be expressed in a single unified word, and there is a preference for comparing the semantic consistency of answers
E Others / Other cases not categorized by the above categories
Appendix H Correlation Experimental Details
Our correlation experiments consist of two parts: we conducted both Reinforcement Learning (RL) (Section H.1) and Rejection Sampling Fine-tuning (RFT) (Section H.2) to validate the usability of VerifyBench.
H.1 Reinforcement Learning Experiments
H.1.1 Setup
Training.
For RL usage, we adopted Qwen2.5-3B-Instruct (Qwen et al., 2025) as the base model and conducted training with the DeepMath-103K (He et al., 2025) dataset. For the RL algorithm, we employed GRPO (Shao et al., 2024) optimization, setting the global batch size to 512 and the mini-batch size to 128. For each query, we sampled 16 responses to compute the group relative advantage. The maximum prompt length was set to 1024, and the maximum response length was set to 7168. For rollouts, we utilized the vLLM (Kwon et al., 2023) engine with temperature fixed at 1.0 and top_p at 1.0. Training was performed with the veRL (Sheng et al., 2025) framework using FSDP as the backend. The learning rate during training was set to 1e-6, with 20 warmup steps. For the LLM verifier, we deployed services through SGLang (Zheng et al., 2024) and performed verification via remote calls. All RL experiments were conducted on 8 H200 GPUs.
Evaluation.
For evaluation, we leveraged the functionality of the veRL framework to conduct assessments on checkpoints every 10 training steps during RL training. To reduce evaluation variance, we set the temperature to 0.7 and top_p to 0.95, sampling each test example 8 times and then computing the average accuracy.
H.1.2 Experimental Results
We applied a sliding-window smoothing method to the curves presented in the main text. In Figure 9, we show the original curves before smoothing (light-colored curves).
Figure 9: The performance(%) of RL across different LLM judges which have various performance on VerifyBench.
H.2 Rejection Sampling Fine-tuning Experiments
H.2.1 Setup